OpenStudy (anonymous):

Solve y = log(3x+3) for x

OpenStudy (jdoe0001):

assuming you've covered logarithms already?

OpenStudy (anonymous):

Applu exponential function on both sides to get rid of log. \[e^y = e^{\log(3x+3)} => e^y = 3x + 3\] work from there.

OpenStudy (anonymous):

3x = e^y - 3